OpenStudy (anonymous):

A geiger-muller tube is a type of gas-filled radiation detector. It can detect particles like X-rays, alpha particles and beta rays (electrons). This is useful in quantizing the activity of a radioactive source or determining if an area containing radioactive material is safe to enter. Assuming that you have one mole of gas, if a geiger counter is filled with 9.330 × 103 Pa of argon gas at room temperature (T = 22.2 C), what is the density of the gas in this geiger tube?

OpenStudy (aaronq):

density is a quantity of mass in a specified volume: \(\rho =\dfrac{m}{V}\) rearrange the ideal gas law, PV=nRT \(PV=nRT=\dfrac{m}{M}RT \rightarrow \dfrac{PM}{RT}=\dfrac{m}{V}=\rho\) \(\large \rho=\dfrac{PM}{RT}\)